OpenStudy (anonymous):

What is the slope of the line that passes through the points (-4, 6) and (2, 4)?

OpenStudy (anonymous):

\[Slope = y _{2} - y _{1}/x _{2} - x _{1}\] or \[y _{1} - y _{2}/ x _{1} - x _{2}\] Using the first formula slope = 4- 6/2-(-4) = -2/6 = -1/3

OpenStudy (anonymous):

or if u wanna use the second one 6-4/-4-2 = 2/-6 = -1/3 Same, same...